# Infinite Dimensional Topological-Holomorphic Symmetry in Three-Dimensions

### v2

@@ 1. Introduction
This is precisely where the three-dimensional topological-holomorphic field theory constructed in [19] enters the picture. It provides a concrete realisation of a quantum field theory with symmetry currents satisfying, schematically, a conservation law of the form $\mathrm{d}^{\prime}J=0$. In [19], it was already anticipated that the symmetries of the theory should form an infinite-dimensional algebra. What was missing, however, was a precise mathematical description of this structure. By recognizing that the differential constraint obeyed by the symmetry currents coincides with the one underlying the raviolo complex of [22], we show that this framework provides exactly the missing ingredient which allows for the explicit construction of the infinite-dimensional symmetry algebra.
@@
In Section 2, we review the Wess–Zumino–Witten (WZW) model and show how the symmetries of this theory give rise, through radial quantisation, to two copies of the Kac–Moody algebra. In Section 3, we introduce the three-dimensional topological-holomorphic field theory constructed in [19], and describe its symmetry structure. In Section 4, we present the raviolo formalism developed in [22] and adapt it to the three-dimensional setting relevant to our theory. In Section 5, we perform the radial quantisation of the three-dimensional theory and compute the commutation relations between the modes of the symmetry currents. In Section 6, we show that these commutation relations define a centrally extended affine graded Lie algebra, which we construct in detail. Finally, in Section 7 we construct the raviolo vertex algebra corresponding to the three-dimensional theory.
@@ 2.1. The Kac-Moody Algebra
where we have used Stokes’ theorem (counterclockwise orientation). Next, since $J(z)=kg^{-1}\partial_{z}g$ is $\mathfrak{g}$–valued, we fix a basis $\{t_{a}\}$ with $\operatorname{Tr}(t_{a}t_{b})=2\delta_{ab}$ and $t_{a},t_{b}]=f_{ab}^{c}t_{c}$, and write
@@
Hence, taking the product of fields to be $X=J^{b}(w)X^{\prime}$, with $X^{\prime}=\prod_{i}\mathcal{O}_{i}(u_{i})$ where $w\in D$ and all $u_{i}\notin D$, and where the $\mathcal{O}_{i}$ denote arbitrary operator insertions, the holomorphy of $\tilde{\alpha}$} on $D$ (extended by zero outside) ensures that $\delta_{\tilde{\alpha}}$ acts only on $J^{b}(w)$. The Ward identity then yields
@@
If we want this relation to hold as an operator identity, we must allow for an arbitrary number of additional fields to appear alongside $a(z)$ and $b(w)$ in a general correlator. In particular, the decomposition of the contour into $c^{1}$ and $C_{2}$ is justified only when $b(w)$ is the unique insertion with a singular OPE with $a(z)$ lying between the two contours.
@@
Let $\mathfrak{G}=(\mathfrak{g}\xrightarrow{\mu_{1}}\mathfrak{h},\mu_{2})$ denote a Lie 2-algebra and let $\mathbb{G}=(\mathsf{H}\xrightarrow{\mathsf{t}}G,\rhd)$ the corresponding Lie 2-group. We consider a three-dimensional manifold $M$, which we take for now to be $M=\mathbb{R}\times\mathbb{C}$. The three-dimensional action is given by [19]
@@
where $g\in C^{\infty}(M,G)$ is a $G$-valued smooth field, $\Theta\in\Omega^{1}(M,\mathfrak{h})$ an $\mathfrak{h}$-valued $1$-form, and $\langle\cdot,\cdot\rangle:\mathfrak{g}\times\mathfrak{h}\to\mathbb{R}$ a degree 1-1, non degenerate, invariant bilinear form. For an introduction to Lie $2$-groups and Lie $2$-algebras, we refer the reader to [34, 35, 36].
@@
Let us pause to highlight a few important observations. First, just like the fields in the theory, the symmetry transformations also come in pairs, once again reflecting the underlying categorical structure. Second, the left-acting symmetry pair is anti-holomorphic, exactly as in the WZW model. In contrast, the right-acting symmetry satisfies a more intricate condition: it involves derivatives with respect to $\bar{z}$ and $\tau$, along with a nontrivial appearance of the map $\mu_{1}$. This marks an important difference from the two-dimensional case. In particular, since the spacetime is now odd-dimensional, there is no natural chiral–anti-chiral decomposition, and we must carefully identify which of these symmetries plays the role of the chiral symmetry familiar from two dimensions—that is, which one gives rise to the infinite-dimensional symmetry algebra.
@@ 3.2. Transverse Holomorphic Foliation
A field theory is said to be topological-holomorphic if its fields are defined on a smooth manifold $M$ equipped with a THF structure $\mathcal{F}$. It is clear that the kinematical and dynamical properties of a topological-holomorphic field theory depend crucially on the underlying THF structure. As we shall see in the upcoming section, the raviolo construction is done for general 3-manifolds which admit a THF.
@@ 4.1. The mixed de Rham-Dolbeault complexes
into the holomorphic Dolbeault differential $\partial$ plus the differential $\mathrm{d}$$^{\prime}$ on the so-called $\mathcal{F}$-basic holomorpic forms [^5] More precisely, these are sheaves of holomorphic functions $f$ on $M$ such that $L_{X}f=0$ for all $X\in V_{F}$.. In this decomposition $V_{\mathcal{F}}^{*}$ is locally spanned by ${\mathrm{d}}\bar{z},{\mathrm{d}}\tau$, while $\big(Q^{(1,0)}\big)^{*}$ is locally spanned by ${\mathrm{d}}z$. Note, for reference, that $\mathcal{A}^{p,q}$ is what we usually denote by $\Omega^{(p,q)}(M)$. Since the space $\mathcal{A}^{p,q}$ will play an important role, let us provide explicit examples of its elements for the lowest degrees.
@@ 4.3. Mode expansion of the currents
Unlike in the Dolbeault complex, no negative powers of $z$ appear here. Consequently, any element $\tilde{\alpha}\in H^{(0,0)}_{\mathrm{d}^{\prime}}(M)$ can be expanded in modes as
@@
(4.3.7) $\Omega^{m}=\frac{(-1)^{m}}{m!}\partial_{z}^{m}\omega\,.$
@@
In other words, the raviolo differential forms serve a role analogous to that of negative modes in ${\mathbb{R}}\times{\mathbb{C}}\setminus\{0\}$, but in the form of 1-forms with non-trivial degree. This is consistent with the expectation from Hartog’s theorem.
@@ 5. Radial Quantisation
Before turning to the details, let us indicate why a very similar construction extends so smoothly to $\mathbb{R}$$\times{\mathbb{C}}$. In the complex plane two ingredients were essential: the mode expansion and Cauchy’s residue theorem, which allowed us to translate the singular behaviour of operator products into commutators. In the three-dimensional setting we have already established the mode expansion of the currents in (4.3.19), and we also introduced the $(1,0)$-form $\omega$ in (4.3.5), which satisfies
@@
This provides a higher–dimensional analogue of Cauchy’s residue formula, allowing us to generalise statements involving circles $S^{1}\subset{\mathbb{C}}$ to spheres $S^{2}\subset{\mathbb{R}}\times{\mathbb{C}}$. In Euclidean signature one can again take the radial coordinate as time and, in direct analogy with the two–dimensional case, define a radial quantisation scheme. The generalised Cauchy formula then makes it possible to translate singular OPEs of the three–dimensional currents into commutation relations.
@@ 5.1. The Ward Identity
Note that the bilinear form $\langle\cdot,\cdot\rangle:\mathfrak{n}\otimes V\oplus V\otimes\mathfrak{n}\to\mathbb{C}$ pairs elements of $\mathfrak{n}$ with those of $V$, and thus, unlike the standard case, it does not provide an identification between either of these spaces and their respective duals. As a result, we cannot use $\kappa$ to raise or lower indices; this is why we define the dual structure constants $\tilde{f}^{ab}_{c}$ and $(\tilde{\mu}_{2})_{c}^{ab}$.
@@ 5.2. Commutation Relations
As we mentioned at the beginning of this section, the essence of this construction is that in analogy to the two-dimensional case, we can also perform radial quantisation. We choose our time coordinate to be the radial coordinate on $\mathbb{R}\times{\mathbb{C}}\cong{\mathbb{R}}^{3}$ so that equal-time slices will be given by spheres of fixed radius. Correlation functionas are then identified as vacuum expectation values of radially ordered products of operators
@@
In §4.3 of  [22], the authors construct a raviolo current algebra by tensoring a Lie superalgebra with the cohomology of the raviolo, and endow it with a bilinear form of even degree. In that setting, the vanishing brackets arise between generators valued in the odd part $\mathfrak{g}_{1}$ of the Lie superalgebra. In contrast, our construction uses a Lie 2-algebra tensored with the same cohomology, but equipped with a bilinear form of degree 1-1. As a consequence, the vanishing brackets in our case occur between degree 0 elements (the $L_{n}$’s), while the non-trivial brackets involve degree 1-1 elements (the $H_{n}$’s). This shift in grading structure can be traced back to the differing degrees of the bilinear forms used in each construction.
Remark 5:
@@ 6.1. The Affine Graded Lie Algebra
as in equation (5.1.13). We assign degree 0 to $V^{\vee}$ and degree 1-1 to $\mathfrak{n}^{\vee}$. With these conventions, we have the following:
@@
The brackets defined in (6.1.4) satisfy the graded Jacobi identity. Hence, they endow $\mathfrak{G}^{\vee}$ with the structure of a graded Lie algebra.
@@
Now denote by ${\mathcal{K}}_{u}$ the raviolo polynomials (4.3.12) locally around the holomorphic variable $u=0$, such that its graded compoents are given by ${\mathcal{K}}_{u}^{0}={\mathbb{C}}[[u]]$ and ${\mathcal{K}}_{u}^{1}=\mathrm{span}_{\mathbb{C}}\{\Omega^{0}_{u},\Omega^{1}_{u},\dots\}$. With this at hand, we define the higher affine graded Lie algebra $\widehat{\mathfrak{G}}$ as
(6.1.9) $\widehat{\mathfrak{G}}=\left(V^{\vee}\otimes{\mathcal{K}}_{u}^{0}\right)\oplus\left(\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}\right).$
The bracket on $\widehat{\mathfrak{G}}$ is defined as follows: For $X\otimes f(u),Y\otimes g(u)\in V^{\vee}\otimes{\mathcal{K}}_{u}^{0}$, we set
@@
(6.1.11) $X\otimes f(u),A\otimes k(\Omega_{u})]=\mu_{2}^{\vee}(X,A)\otimes f(u)k(\Omega_{u})\in\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}\,.$
(6.1.12) $A\otimes k(\Omega_{u}),B\otimes l(\Omega_{u})]=0\,.$
@@
With this definition at hand, we can show that the graded Lie brackets reproduce the commutators of the higher currents, up to the central term. Specifically, consider the bijective graded linear map
@@
which identifies the modes of the currents with the generators of the affine graded Lie algebra. Using the relations in equation (6.1.13), one readily verifies that the non-central contributions to the commutators in equations (5.2.7) and (5.2.8) coincide precisely with the graded Lie bracket on $\widehat{\mathfrak{G}}$.
@@ 6.2. The Central Extension
(6.2.1) $\frac{1}{8\pi i}\oint_{S^{2}}\mathrm{d}z\wedge u^{n}\Omega^{m}_{u}=\delta_{n,m}\,.$
@@
The notation ${\mathbb{C}}[-1]$ reflects the fact that we are working with chain complexes. Concretely, it indicates that $\operatorname{Res}$ vanishes on $(0,0)$-forms, while its action on $(1,0)$-forms yields values in $\mathbb{C}$}, as explicitly shown in (6.2.2).
@@
We aim to show that the bilinear pairing $\langle\cdot,\cdot\rangle^{\vee}$ on $\widehat{\mathfrak{G}}$ is invariant under the dual Lie bracket structure, specifically that
@@
The bilinear map $\mathfrak{w}:\widehat{\mathfrak{G}}\times\widehat{\mathfrak{G}}\to{\mathbb{C}}$ determines a graded Lie algebra 2-cocycle.
@@
The Poisson algebra of the currents of the 3d theory (4.2.1) defines a centrally extended affine Lie algebra $\widehat{\mathfrak{G}}_{\mathfrak{w}}$ fitting into the central extension sequence
(6.2.16) $0\rightarrow{\mathbb{C}}[1]\rightarrow\widehat{\mathfrak{G}}_{\mathfrak{w}}\rightarrow\widehat{\mathfrak{G}}\rightarrow 0\,$,

### v3

@@ 1. Introduction
This is precisely where the three-dimensional field theory constructed in [19] enters the picture. It provides a concrete realisation of a quantum field theory with symmetry currents satisfying, schematically, a conservation law of the form $\mathrm{d}^{\prime}J=0$. In [19], it was already anticipated that the symmetries of the theory should form an infinite-dimensional algebra. What was missing, however, was a precise mathematical description of this structure. By recognizing that the differential constraint obeyed by the symmetry currents coincides with the one underlying the raviolo complex of [22], we show that this framework provides exactly the missing ingredient which allows for the explicit construction of the infinite-dimensional symmetry algebra.
@@
In Section 2, we review the Wess–Zumino–Witten (WZW) model and show how the symmetries of this theory give rise, through radial quantisation, to two copies of the Kac–Moody algebra. In Section 3, we introduce the three-dimensional theory constructed in [19], and describe its symmetry structure. In Section 4, we present the raviolo formalism developed in [22] and adapt it to the three-dimensional setting relevant to our theory. In Section 5, we perform the radial quantisation of the three-dimensional theory and compute the commutation relations between the modes of the symmetry currents. In Section 6, we show that these commutation relations define a centrally extended affine graded Lie algebra, which we construct in detail. Finally, in Section 7 we construct the raviolo vertex algebra corresponding to the three-dimensional theory.
@@ 2.1. The Kac-Moody Algebra
where we have used Stokes’ theorem (with counterclockwise orientation). Next, since $J(z)=kg^{-1}\partial_{z}g$ is $\mathfrak{g}$–valued, we fix a basis $\{t_{a}\}$ with $\operatorname{Tr}(t_{a}t_{b})=2\delta_{ab}$ and $t_{a},t_{b}]=f_{ab}^{c}t_{c}$ in order to write
@@
Hence, taking the product of fields to be $X=J^{b}(w)X^{\prime}$ where $X^{\prime}=\prod_{i}\mathcal{O}_{i}(u_{i})$ is a product of arbitrary operator insertions with $w\in D$ and $u_{i}\notin D$, the holomorphy of $\tilde{\alpha}$} on $D$ (extended by zero outside) ensures that $\delta_{\tilde{\alpha}}$ acts only on $J^{b}(w)$. The Ward identity then yields
@@
If we want this relation to hold as an operator identity, we must allow for an arbitrary number of additional fields to appear alongside $a(z)$ and $b(w)$ in a general correlator. In particular, the decomposition of the contour into $C_{1}$ and $C_{2}$ is justified only when $b(w)$ is the unique insertion with a singular OPE with $a(z)$ lying between the two contours.
@@
Let $\mathfrak{G}=(\mathfrak{g}\xrightarrow{\mu_{1}}\mathfrak{h},\mu_{2})$ denote a (real) Lie 2-algebra and let $\mathbb{G}=(\mathsf{H}\xrightarrow{\mathsf{t}}G,\rhd)$ the corresponding Lie 2-group. We consider a three-dimensional manifold $M$, which we take for now to be $M=\mathbb{R}\times\mathbb{C}$. The three-dimensional action is given by [19]
@@
where $g\in C^{\infty}(M,G)$ is a $G$-valued smooth field, $\Theta\in\Omega^{1}(M,\mathfrak{h})$ an $\mathfrak{h}$-valued $1$-form, and $\langle\cdot,\cdot\rangle:\mathfrak{g}\times\mathfrak{h}\to\mathbb{C}$ a degree 1-1, non degenerate, invariant bilinear form. For an introduction to Lie $2$-groups and Lie $2$-algebras, we refer the reader to [34, 35, 36].
@@
Let us pause to highlight a few important observations. First, just like the fields in the theory, each left-/right-acting symmetry transformation parameter comes in a pair, once again reflecting the underlying categorical structure. Second, the left-acting symmetry pair is anti-holomorphic, exactly as in the WZW model. In contrast, the right-acting symmetry satisfies a more intricate condition: it involves derivatives with respect to $\bar{z}$ and $\tau$, along with a nontrivial appearance of the differential $\mu_{1}$. This marks an important difference from the two-dimensional case. In particular, since the spacetime is now odd-dimensional, there is no natural chiral–anti-chiral decomposition, and we must carefully identify which of these symmetries plays the role of the chiral symmetry familiar from two dimensions—that is, which one gives rise to the infinite-dimensional symmetry algebra.
@@ 4.1. The mixed de Rham-Dolbeault complexes
into the holomorphic Dolbeault differential $\partial$ plus the differential $\mathrm{d}$$^{\prime}$ on the so-called $\mathcal{F}$-basic holomorpic forms [^5] More precisely, these are sheaves of holomorphic functions $f$ on $M$ such that the Lie derivative $L_{X}f=0$ vanishes for all $X\in V_{F}$.. In this decomposition $V_{\mathcal{F}}^{*}$ is locally spanned by ${\mathrm{d}}\bar{z},{\mathrm{d}}\tau$, while $\big(Q^{(1,0)}\big)^{*}$ is locally spanned by ${\mathrm{d}}z$. Note, for reference, that $\mathcal{A}^{p,q}$ is what we usually denote by $\Omega^{(p,q)}(M)$. Since the space $\mathcal{A}^{p,q}$ will play an important role, let us provide explicit examples of its elements for the lowest degrees.
@@ 4.3. Mode expansion of the currents
Unlike in the Dolbeault complex, no negative powers of $z$ appear here; this is a manifestation of Hartogs’s theorem (see also [41]). Consequently, any element $\tilde{\alpha}\in H^{(0,0)}_{\mathrm{d}^{\prime}}(M)$ can be expanded in modes as
@@
(4.3.7) $\Omega^{m}=\frac{(-1)^{m}}{m!}\partial_{z}^{m}\omega\,,\qquad\Omega^{0}=\omega\,.$
@@
In other words, the raviolo differential forms serve a role analogous to that of negative modes in ${\mathbb{R}}\times{\mathbb{C}}\setminus\{0\}$, but in the form of 1-forms with non-trivial degree.
@@ 5. Radial Quantisation
Before turning to the details, let us indicate why a very similar construction extends so smoothly to $\mathbb{R}$$\times{\mathbb{C}}$. In the complex plane, two ingredients were essential: the mode expansion and Cauchy’s residue theorem, which allowed us to translate the singular behaviour of operator products into commutators. In the three-dimensional setting, we have already established the mode expansion of the currents in (4.3.19), and we also introduced the Bochner-Martinelli $(1,0)$-form $\omega$ in (4.3.5), which satisfies
@@
This provides a higher–dimensional analogue of Cauchy’s residue formula, allowing us to generalise statements involving circles $S^{1}\subset{\mathbb{C}}$ to spheres $S^{2}\subset{\mathbb{R}}\times{\mathbb{C}}$. In Euclidean signature, one can again take the radial coordinate as time and, in direct analogy with the two–dimensional case, define a radial quantisation scheme. The generalised Cauchy formula then makes it possible to translate singular OPEs of the three–dimensional currents into commutation relations.
@@ 5.1. The Ward Identity
Note that the bilinear form $\langle\cdot,\cdot\rangle:\mathfrak{n}\times V\to\mathbb{C}$ pairs elements of $\mathfrak{n}$ with those of $V$, and thus, unlike the standard case, it does not provide an identification between either of these spaces and their respective duals. As a result, we cannot use $\kappa$ to raise or lower indices; this is why we define the dual structure constants $\tilde{f}^{ab}_{c}$ and $(\tilde{\mu}_{2})_{c}^{ab}$.
@@ 5.2. Commutation Relations
As we mentioned at the beginning of this section, the essence of this construction is that in analogy to the two-dimensional case, we can also perform radial quantisation. We choose our time coordinate to be the radial coordinate on $\mathbb{R}\times{\mathbb{C}}\cong{\mathbb{R}}^{3}$ so that equal-time slices will be given by spheres of fixed radius. Correlation functions are then identified as vacuum expectation values of radially ordered products of operators
@@ 6.1. The Affine Graded Lie Algebra
as in equation (5.1.13). In accordance with the literature [44, 45, 46, 47], we will assign the cohomological degree -1 to $\mathfrak{n}^{\vee}$, and cohomological degree 0 to $V^{\vee}$. The reason for this is given in Remark 5.
Remark 5:
Given a Lie 2-algebra $\mathfrak{G}=\mathfrak{h}\xrightarrow{\mu_{1}}\mathfrak{g}$, the dual Lie 2-algebra $\mathfrak{G}^{\vee}[1]=\mathfrak{g}^{\vee}\xrightarrow{\mu_{1}^{\vee}}\mathfrak{h}^{\vee}$ is given by the underlying dual complex together with a shift in degree down by $1$, denoted by ”[1][1]”, such that $\mathfrak{g}^{\vee}$ has degree 1-1 and $\mathfrak{h}^{\vee}$ has degree 0. This is to ensure that $\mu_{1}^{\vee}$ remains a cohomological differential which increases the degree by $1$, and hence $\mathfrak{G}^{\vee}[1]$ can still be treated as a Lie 2-algebra.
With the above grading conventions, we have the following:
@@
The brackets defined in (6.1.4) satisfy the graded Jacobi identity. Hence, they endow $\mathfrak{G}^{\vee}=\mathfrak{G}^{\vee}[-1]$ with the structure of a graded Lie algebra.
@@
We denote by ${\mathcal{K}}_{u}^{\bullet}$ the formal raviolo power series [^10] Here, $\mathcal{K}$ is the version of the raviolo polynomials $\mathcal{K}_{\text{poly}}$ (4.3.12) with $\mathbb{C}[u]$ in degree 0 replaced by the formal power series $\mathbb{C}[\![u]\!]$. locally around the holomorphic variable $u=0$, such that its graded components are given by ${\mathcal{K}}_{u}^{0}={\mathbb{C}}[\![u]\!]$ and ${\mathcal{K}}_{u}^{1}=\mathrm{span}_{\mathbb{C}}\{\Omega^{0}_{u},\Omega^{1}_{u},\dots\}$. With this at hand, we define the higher affine graded Lie algebra $\widehat{\mathfrak{G}}^{\vee}$ as the total degree 0 piece of the tensor product complex
(6.1.9) $\widehat{\mathfrak{G}}^{\vee}=\big(\mathfrak{G}^{\vee}\otimes\mathcal{K}^{\bullet}_{u}\big)_{0}=\left(V^{\vee}\otimes{\mathcal{K}}_{u}^{0}\right)\oplus\left(\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}\right)\,.$
The bracket on $\widehat{\mathfrak{G}}^{\vee}$ is defined as follows: For $X\otimes f(u),Y\otimes g(u)\in V^{\vee}\otimes{\mathcal{K}}_{u}^{0}$, we set
@@
For $X\otimes f(u)\in V^{\vee}\otimes{\mathcal{K}}_{u}^{0}$ and $A\otimes k(u)\in\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}$, the bracket is
(6.1.11) $X\otimes f(u),A\otimes k(u)]=\mu_{2}^{\vee}(X,A)\otimes f(u)k(\Omega_{u})\in\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}\,.$
Finally, for two elements $A\otimes k(u),B\otimes l(u)\in\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}$, the bracket vanishes:
(6.1.12) $A\otimes k(u),B\otimes l(u)]=0\,.$
@@
In accordance with the above bracket structure, we shall assign a homological grading to $\widehat{\mathfrak{G}}^{\vee}$ such that
(6.1.14) $\text{deg}\left(V^{\vee}\otimes{\mathcal{K}}_{u}^{0}\right)=0,\qquad\text{deg}\left(\mathfrak{n}^{\vee}\otimes{\mathcal{K}}_{u}^{1}\right)=1\,.$
We shall adopt this degree convention in the rest of the following.
We can show that the graded Lie brackets reproduce the commutators of the higher currents, up to the central term. Specifically, consider the bijective graded linear map
@@
which identifies the modes of the currents with the generators of the affine graded Lie algebra. Using the relations in equation (6.1.13), one readily verifies that the non-central contributions to the commutators in equations (5.2.7) and (5.2.8) coincide precisely with the graded Lie bracket on $\widehat{\mathfrak{G}}^{\vee}$.
@@ 6.2. The Central Extension
(6.2.1) $\frac{1}{8\pi i}\oint_{S^{2}}\mathrm{d}u\wedge u^{n}\Omega^{m}_{u}=\delta_{n,m}\,.$
@@
The notation ${\mathbb{C}}[-1]$ reflects the fact that we are working with a chain complex with $\mathbb{C}$} in degree 1 — namely, the notation ”[1][-1]” indicates that the degree has been shifted up by 1; cf. Remark 5. Concretely, it indicates that $\operatorname{Res}$ vanishes on $(0,0)$-forms, while its action on $(1,0)$-forms yields values in $\mathbb{C}$}, as explicitly shown in (6.2.2).
@@
We aim to show that the bilinear pairing $\langle\cdot,\cdot\rangle^{\vee}$ on $\mathfrak{G}^{\vee}$ is invariant under the dual Lie bracket structure, specifically that
@@
The bilinear map $\mathfrak{w}:\widehat{\mathfrak{G}}^{\vee}\times\widehat{\mathfrak{G}}^{\vee}\to{\mathbb{C}}[-1]$ determines a graded Lie algebra 2-cocycle.
@@
The Poisson algebra of the currents of the 3d theory (4.2.1) defines a centrally extended affine Lie algebra $\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}$ fitting into the central extension sequence
(6.2.16) $0\rightarrow{\mathbb{C}}[-1]\rightarrow\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}\rightarrow\widehat{\mathfrak{G}}^{\vee}\rightarrow 0\,$,